Which property can help you to mentally evaluate 26(19)?
Answer: Communative Property of Multipication

Answer by jim_thompson5910(35256)

19 = 20-1

So

26(19) = 26(20-1) = 26*20 - 26*1 = 520-26 = 494

I'm using the distributive property when going from 26(20-1) to 26*20 - 26*1